The cart and package have a mass of 20 kg and 5 kg respectively. If the card has a smooth surface and it is initially at rest, while the velocity of the package is as shown, determine the final velocity of the cart and package after in the impact.


1
Expert's answer
2021-05-27T08:31:31-0400

"m_c=20\;kg \\

m_p = 5\;kg \\

(v_p)_1=10\;m/s"

The expression of the velocity in the horizontal direction:

"(v_p)_{1H}=\frac{4}{5}(v_p)_1 \\

(v_p)_{1H}=\frac{4}{5} \times 10 = 8 \;m/s"

Since the external force is zero, therefore the momentum is conserved. Write the expression of the conservation of the momentum.

"m_c(v_c)_{1H}+m_p(v_p)_{1H}=m_c(v_2)+m_p(v_2) \\

m_c(v_c)_{1H}+m_p(v_p)_{1H}=(m_c+m_p)(v_2) \\

20 \times 0 + 5 \times 8 = (20 + 5)(v_2) \\

v_2 = \frac{40}{25}=1.6 \;m/s"

Answer: 1.6 m/s